# Joint estimation of a two-phase spin rotation beyond classical limit

### v1

@@ Abstract
The gray dots show the measured squeezing parameter when only one spin-nematic subspace is detected using the measurement procedure identical to the earlier study [31]. The blue and orange dots illustrate quantum noise reduction in both SU(2) subspaces spanned by $\left\{L_{y},Q_{xz},\left(Q_{xx}-Q_{zz}\right)/2\right\}$ and $\left\{Q_{yz},L_{x},\left(Q_{yy}-Q_{zz}\right)/2\right\}$. The squeezing parameters for the above two should ideally be equal in theory, however, imperfections of the synthesized MW pulses before atom number detection bring in additional noise such as from the fluctuations of signal ($a_{S}$ and $a_{A}$) amplitude and pump ($a_{P}$) phase (see Supplementary information), which degrades the detected squeezing. For longer evolution times when the system reaches the over-squeezed regime, spin-nematic squeezing parameters measured in both subspaces are found lower due to the breakdown of undepleted pump approximation. The observed squeezing for both $Q_{xz}$ and $Q_{yz}$ indirectly benchmark our capability of synthesizing the above MW pulse sequence shown in Fig. 1(c).
@@
The phase encoding operation $e^{i\left(\varphi_{1}L_{x}+\varphi_{2}L_{y}\right)}$ is enacted by small angle RF rotations, which takes about 100 $\mu$s to complete. To ensure maximum sensitivity of the probe state to the phase imprinting procedure, the rotation axis is aligned with the anti-squeezed quadrature as depicted in Fig. 2(a). In the alternatively parameterized form $\varphi_{1}=\alpha\cos\Theta$ and $\varphi_{2}=\alpha\sin\Theta$ as shown in Fig. 2(b), $\alpha$ of the imprinted phase is controlled experimentally by the duration of the RF pulses while $\Theta$ by the RF phases.
@@
Figure 3: Metrological gains for joint estimation of two spin rotations. The Metrological gain obtained from Eq. (4) with error propagation formulae $\Delta\varphi_{1}=\Delta Q_{yz}/\left|d\left\langle Q_{yz}\right\rangle/d\varphi_{1}\right|$ and $\Delta\varphi_{2}=\Delta Q_{xz}/\left|d\left\langle Q_{xz}\right\rangle/d\varphi_{2}\right|$, where the denominators and numerators are obtained by fitting the measured spin-nematic operators $\langle Q_{xz(yz)}^{\mathrm{meas}}\rangle$ and $\Delta Q_{xz(yz)}^{\mathrm{meas}}$ using quartic functions without linear terms, respectively. Gains of $3.3$ dB to $6.3$ dB over the SQL are observed for $\varphi_{1,2}\in[-0.05,0.05]\pi$.
@@
Figure 2(c-f) show the parameter-dependent mean values and standard deviations of the measured $Q_{yz}$ and $Q_{xz}$. Using the non-entangled polar state as the probe, the standard deviation observed is slightly higher than the shot noise as a result of technical imperfections of the pulses. For the spin-nematic squeezed state, the distributions of measured data are found squeezed along both the $Q_{xz}$ and $Q_{yz}$ axes. The mean values are linearly dependent on the parameter $\alpha$ for fixed $\Theta$, while the standard deviations are almost independent of the parameters $\alpha$ and $\Theta$. The measurement microwave pulses are generated by the same AWG (arbitrary waveform generator) as the phase imprinting RF pulses, which ensures a nearly perfect phase correlation between the two procedures. The total read-out protocol takes about 374 $\mu$s, and during which the less-than-ideal differential phase shift between the two microwave signals also contributes to the phase noise, which altogether limits the measured squeezing and therefore the realized metrological gain. To further elucidate the performance of the joint estimation scheme, in Fig. 3, the metrological gain for the two-phase estimation, defined as
@@
is shown in the rotation range of $\varphi_{1,2}\in[-0.05,0.05]\pi$ encoded to the spin-nematic squeezed probe state, enhancement of $3.3$ dB to $6.3$ dB beyond the SQL of 26000 atoms is observed. About 500 atoms are lost during the experiment cycle. We have also conducted an analysis of the metrological gain of the polar state and a sensitivity approximately 2 dB worse than the SQL is found, due to the crosstalk noise (See supplementary Material). The metrological gain is expected to be symmetric with respect to $\varphi_{1}$ and $\varphi_{2}$, while the experimental results are affected by the pump phase noise (see Supplementary information) and degrade stochastically in the $\varphi_{1}$-$\varphi_{2}$ plane. Such degradation of metrological gain can be mitigated to some degree by carefully tuning the pump phase before sensing.
In conclusion, we propose and experimentally demonstrate, for the first time, the simultaneous estimation of two phases generated by two orthogonal (noncommuting) spin rotations using an entangled atomic ensemble. Sensitivities exceeding 3 dB beyond the standard quantum limit are observed across a wide range of rotation angles. By coherently coupling the atomic BEC in the ground $F=1$ manifold to the $F=2$ manifold, with the latter serving as the auxiliary field, and using a carefully designed microwave pulse sequence, simultaneous readout of two noncommuting observables is realized. These observables serve as estimators for the broad range of parameters being measured. The metrological advantage of spin-nematic squeezed state was previously demonstrated in single-parameter estimation scenarios [41]. This work expands such quantum-enhanced sensing applications to multi-parameter quantum metrology and will likely stimulate the development of a broad range of application scenarios in atomic sensors. Limited by the coherence time of the spin-nematic squeezed state in our experiment, interaction-based readout, which recently enabled significantly enhanced metrological gains over direct linear read-out adopted here [41], remains to be implemented for studying the two-parameter estimation of this work. An operational protocol is known [33] and whose implementation will mitigate a large part of the atom number detection noise after coherence time is improved.
We thank Drs Q. Liu and L. N. Wu for helpful discussions. This work is supported by the National Key R&D Program of China (Grants No. 2018YFA0306504), by the National Natural Science Foundation of China (NSFC) (Grants No. 92265205 and No. U1930201), and by the Innovation Program for Quantum Science and Technology (2021ZD0302100).
@@
I Spin-nematic squeezing in spin-1 system
@@ I.1 Dynamical simulations
For a spin-1 Bose-Einstein condensate (BEC) composed of $N$ ${}^{87}$b atoms, the full spin-mixing dynamics under single spatial mode approximation (SMA) [36] is described by the following Hamiltonian
$H=$ $\frac{c_{2}}{2N}\bigl{[}2(a_{1}^{\dagger}a_{-1}^{\dagger}a_{0}a_{0}+a_{1}a_{-1}a_{0}^{\dagger}a_{0}^{\dagger})+(N_{1}-N_{-1})^{2}$ (S1)
$+(2N_{0}-1)(N_{1}+N_{-1})\big{]}-qN_{0},$
@@
where $\mathcal{D}[a_{m}]\rho=a_{m}\rho a_{m}^{\dagger}-\{a_{m}^{\dagger}a_{m},\rho\}/2$ accounts for atom loss with dissipative rate $\gamma_{m}$ for the $|F=1,m\rangle$ component. Here we assume $\gamma_{m}=\gamma$ for $m=0,\pm 1$, and $H(t)$ is a time-dependent Hamiltonian obtained after replacing $N$ and $c_{2}$ in Eq. (S1) by $N(t)=Ne^{-\gamma t}$ and $c_{2}(t)=c_{2}e^{-\gamma_{c}t}$, where $\gamma_{c}$ is the decay rate of $c_{2}$.
@@
In the above, $c_{2}^{\prime}=c_{2}(t)/N(t)$ denotes the drifting spin-exchange rate in the presence of atom loss. $d\xi_{m}(t)$ represents a complex Wiener noise increment satisfying $\overline{d\xi_{m}(t)}=0$, $\overline{d\xi_{m}^{*}(t)d\xi_{n}(t)}=\delta_{m,n}dt$. The SDEs are capable of describing quantum evolution of the Master Eq. (S2) after averaged over an ensemble of trajectories. The effect of quantum noise is incorporated into the probability distribution of the initial polar state, which is given by the Wigner function and can be sampled according to
@@
where $\alpha_{m}$ and $\beta_{m}$ are independent real numbers following a standard normal distribution. The averages and variances of mode populations are then obtained from the averages of $10000$ simulated trajectories, calculated according to $\langle{\hat{N}}_{m}\rangle=\overline{\psi_{m}^{*}\psi_{m}}-1/2$ and $(\Delta\hat{N}_{m})^{2}=(\Delta\psi_{m}^{*}\psi_{m})^{2}-1/4$.
@@ I.2 SU(3) operators and Bloch spheres
To better understand the spin-mixing dynamics in phase space, both the angular momentum $L_{i}$ and the nematic tensor $Q_{i,j}$ ($i,j\in x,y,z$) operators, which constitute SU(3) Lie algebra, are needed. The definition of the nematic tensor operators are as follows
@@
$Q_{xx}=\frac{2}{3}a_{0}^{\dagger}a_{0}-\frac{1}{3}a_{1}^{\dagger}a_{-1}-\frac{1}{3}a_{-1}^{\dagger}a_{-1}+a_{-1}^{\dagger}a_{1}+a_{1}^{\dagger}a_{-1},$
$Q_{yy}=\frac{2}{3}a_{0}^{\dagger}a_{0}-\frac{1}{3}a_{1}^{\dagger}a_{-1}-\frac{1}{3}a_{-1}^{\dagger}a_{-1}-\hat{a}_{-1}^{\dagger}a_{1}-a_{1}^{\dagger}a_{-1},$
@@ II.1 Quadrature squeezing and atomic homodyne detection
For the short-time evolution starting from polar state, atom number in the $|F=1,m=0\rangle$ component is approximately undepleted. At $q=|c_{2}|$, the complete Hamiltonian approximately takes the two-mode squeezing form after the replacement $\hat{a}_{0}^{\dagger}\rightarrow\sqrt{N}e^{i\phi_{0}}$
@@
with $\phi_{0}=0$. Equation (S7) thus gives the equivalent two-mode squeezing dynamics as in undepleted pump optical four-wave mixing (FWM) or parametric down conversion, and the angular momentum operators $L_{x}$ and $L_{y}$ as well as the nematic tensor operators $Q_{yz}$ and $Q_{xz}$ all reduce to quadrature operators
@@
In the simultaneous measurement scheme adopted for our experiment as depicted in Fig. 1(c) of the main text, the first three steps execute the transformation of
@@
Here for simplicity, we take $\varepsilon_{1}\approx\varepsilon_{2}\approx 1/\sqrt{2}$, while in realistic experiments a small number of atoms (less than $1000$) could remain in the $|F=1,m=0\rangle$ component, which can be neglected as the population of the pump is relatively large.
The fourth step coupling the signals $a_{S}$ and $a_{A}$ to the coherent pump enable atomic homodyne detection simultaneously. After the MW pulse, we obtain
@@
II.2 Influence of crosstalk between the pump and the signal

### v2

@@ Abstract
The gray dots show the measured squeezing parameter when only one spin-nematic subspace is detected [41]. The blue and orange dots illustrate quantum noise reduction in the two spin-nematic SU(2) subspaces. Their respective squeezing parameters ideally should be equal in theory due to the relation $Q_{xz}=e^{iL_{z}\pi/2}Q_{yz}e^{-iL_{z}\pi/2}$, as the spin-nematic state resides within the zero-magnetization subspace $L_{z}=0$, ensuring $\Delta Q_{yz(xz)}=\Delta Q_{yz(xz)}$, or the equality of their variances. However, imperfections of the synthesized MW pulses before atom number detection bring in additional noise, e.g. from the fluctuations of signal ($a_{S}$ and $a_{A}$) amplitude and pump ($a_{2,\pm 1}$) phase (see SM), which together degrades the measured squeezing. For longer evolution times, spin-nematic squeezing parameters measured in both subspaces are found worse due to the breakdown of undepleted pump approximation. Nevertheless, the observed high level of squeezing for both $Q_{xz}$ and $Q_{yz}$ indirectly benchmark our capability of synthesizing the effective transformations shown in Fig. 1(b).
@@
The phase encoding operation $e^{i\left(\varphi_{1}L_{x}+\varphi_{2}L_{y}\right)}$ takes about 100 $\mu$s to complete and is enacted by small angle RF rotations. To ensure maximum sensitivity of the probe state to the phase imprinting procedure, the rotation axis is aligned with the antisqueezed quadrature as depicted in Fig. 2(a). In the alternatively parameterized form $\varphi_{1}=\alpha\cos\Theta$ and $\varphi_{2}=\alpha\sin\Theta$ as shown in Fig. 2(b), $\alpha$ of the imprinted phase is controlled experimentally by the duration of the RF pulse while $\Theta$ by the RF phase.
@@
Figure 3: Metrological gains for joint estimation of two-phase rotations according to Eq. (4) with error propagation formulas $\Delta\varphi_{1}=\Delta Q_{yz}/\left|d\left\langle Q_{yz}\right\rangle/d\varphi_{1}\right|$ and $\Delta\varphi_{2}=\Delta Q_{xz}/\left|d\left\langle Q_{xz}\right\rangle/d\varphi_{2}\right|$ respectively. The denominators and numerators are obtained by fitting the measured spin-nematic operators $\langle Q_{xz(yz)}^{\mathrm{meas}}\rangle$ and $\Delta Q_{xz(yz)}^{\mathrm{meas}}$ using quartic functions without linear terms, respectively. Gains of 3.33.33.3 dB to 6.36.36.3 dB over the SQL are observed for $\varphi_{1,2}\in[-0.05,0.05]\pi$.
@@
Figures 2(c)-(f) show the parameter-dependent mean values and standard deviations of the measured $Q_{yz}$ and $Q_{xz}$ using the nonentangled polar state and the spin-nematic squeezed state as probes. For the polar state, the standard deviation observed is found to be slightly larger than the shot noise as a result of technical imperfections. For the spin-nematic squeezed state, the distributions of measured data are squeezed along both the $Q_{xz}$ and $Q_{yz}$ axes. Their mean values are linearly dependent on the parameter $\alpha$ for fixed $\Theta$, while the standard deviations are almost independent of the parameters $\alpha$ and $\Theta$. The measurement MW pulses are generated by the same AWG as the phase encoding RF pulses to ensure nearly perfect phase correlation between the two steps. The total measurement protocol takes about 374 $\mu$s to complete, during which the less-than-ideal differential phase shift between the two MW signals starts to contribute to phase noise, which altogether limits the measured squeezing and therefore the achieved metrological gain.
To further elucidate the performance of the joint estimation protocol, in Fig. 3, the metrological gain for the two-phase estimation, defined as
@@
is shown in the rotation range of $\varphi_{1,2}\in[-0.05,0.05]\pi$, enhancement of 3.33.33.3 to 6.36.36.3 dB beyond the SQL of 26 000 atoms is observed, despite of the loss of about 500 atoms during the experiment cycle. We have also conducted an analysis of the metrological gain for the polar state and a sensitivity approximately 2 dB worse than the SQL is found, due to the crosstalk noise (see SM). While the metrological gain is expected to be symmetric with respect to $\varphi_{1}$ and $\varphi_{2}$, the experimental results are affected by pump phase noise (see SM) and degrade stochastically in the $\varphi_{1}$-$\varphi_{2}$ plane.
In conclusion, we propose and experimentally demonstrate, for the first time, the simultaneous estimation of two phases generated by two orthogonal spin rotations using an entangled atomic BEC. Coherent coupling of the $F=1$ to the $F=2$ manifold, with the latter serving as an auxiliary field, enables simultaneous measurement of two observables, from which two parameters are estimated. Sensitivities exceeding 3 dB beyond the SQL are observed across a wide range of rotation angles. The metrological advantage of spin-nematic squeezed state was previously demonstrated for single-parameter estimation scenarios [50]. The present work expands such quantum-enhanced sensing applications to multiparameter metrology and will likely stimulate the development of a broad range of applications in atomic sensors. Limited by the coherence time of the spin-nematic squeezed state in our experiment, interaction-based readout, which recently enabled significantly enhanced metrological gains over direct linear read-out adopted here [50], remains to be implemented. An operational protocol is known [51] and its implementation in the future will mitigate a large part of the atom number detection noise after coherence time is improved.
We thank Drs Q. Liu and L. N. Wu for helpful discussions. This work is supported by the National Natural Science Foundation of China (NSFC) (Grants No. 92265205 and No. 12361131576), and by the Innovation Program for Quantum Science and Technology (2021ZD0302100).
@@
I Spin-nematic squeezing in spin-1 BEC
@@ I.1 Dynamical simulations
For a spin-1 Bose-Einstein condensate (BEC) composed of $N$ 87Rb atoms, the full spin-mixing dynamics under single spatial mode approximation (SMA) [45] is described by the following Hamiltonian
$H=$ $\frac{c_{2}}{2N}\bigl{[}2(a_{1}^{\dagger}a_{-1}^{\dagger}a_{0}a_{0}+a_{1}a_{-1}a_{0}^{\dagger}a_{0}^{\dagger})+(N_{1}-N_{-1})^{2}+(2N_{0}-1)(N_{1}+N_{-1})\big{]}-qN_{0},$ (S1)
@@
where $\mathcal{D}[a_{m}]\rho=a_{m}\rho a_{m}^{\dagger}-\{a_{m}^{\dagger}a_{m},\rho\}/2$ accounts for atom loss with dissipative rate $\gamma_{m}$ for the $|F=1,m\rangle$ component. Here we assume $\gamma_{m}=\gamma$ for $m=0,\pm 1$, and $H(t)$ is a time-dependent Hamiltonian obtained after replacing $N$ and $c_{2}$ in Eq. (S1) by $N(t)=Ne^{-\gamma t}$ and $c_{2}(t)=c_{2}e^{-\gamma_{c}t}$, where $\gamma_{c}$ is the measured decay rate of $c_{2}$.
@@
In the above, $c_{2}^{\prime}=c_{2}(t)/N(t)$ denotes the drifting spin-exchange rate in the presence of atom loss. $d\xi_{m}(t)$ represents a complex Wiener noise increment satisfying $\overline{d\xi_{m}(t)}=0$ and $\overline{d\xi_{m}^{*}(t)d\xi_{n}(t)}=\delta_{m,n}dt$. The above SDEs can describe quantum evolution of the Master Eq. (S2) after averaged over an ensemble of trajectories. The effect of quantum noise is incorporated into the probability distribution of the initial polar state, which is given by the Wigner function and can be sampled according to
@@
where $\alpha_{m}$ and $\beta_{m}$ are independent real numbers following a standard normal distribution. The averages and variances of mode populations are then obtained from the averages of simulated trajectories, calculated according to $\langle{\hat{N}}_{m}\rangle=\overline{\psi_{m}^{*}\psi_{m}}-1/2$ and $(\Delta\hat{N}_{m})^{2}=(\Delta\psi_{m}^{*}\psi_{m})^{2}-1/4$. We take 10000 trajectories for most of this study.
@@ I.2 SU(3) operators and Bloch spheres
To better understand the spin-mixing dynamics in phase spaces, both the angular momentum $L_{i}$ and the nematic tensor $Q_{i,j}$ ($i,j\in x,y,z$) operators, which constitute SU(3) Lie algebra, are needed. The definition of the nematic tensor operators are as follows
@@
$Q_{xx}=\frac{2}{3}a_{0}^{\dagger}a_{0}-\frac{1}{3}a_{1}^{\dagger}a_{1}-\frac{1}{3}a_{-1}^{\dagger}a_{-1}+a_{-1}^{\dagger}a_{1}+a_{1}^{\dagger}a_{-1},$
$Q_{yy}=\frac{2}{3}a_{0}^{\dagger}a_{0}-\frac{1}{3}a_{1}^{\dagger}a_{1}-\frac{1}{3}a_{-1}^{\dagger}a_{-1}-\hat{a}_{-1}^{\dagger}a_{1}-a_{1}^{\dagger}a_{-1},$
@@
I.3 Validity of the single-mode approximation in our experiment
In the experiment, we prepare a condensate of $N=26500\pm 250$ 87Rb atoms, tightly confined within an optical dipole trap with frequencies $\omega_{x,y,z}=2\pi\times\left(118,92,185\right)$Hz. The s-wave scattering lengths for the total spin-0 and spin-2 channels are $a_{0}=101.8a_{B}$ and $a_{2}=100.4a_{B}$ ($a_{B}$ denotes the Bohr radius), respectively. Consequently, the Thomas-Fermi radii of the condensate are $R_{x}=4.8\mu$m, $R_{y}=6.1\mu$m and $R_{z}=3.0\mu$m. These parameters yield a spin healing length of approximately $25\mu$m, which is much larger than the size of the condensate. Furthermore, the magnetic field gradient in our setup is sufficiently small, ensuring that the overall trapping potential is the same for all three spin components. Given these considerations, the condensate can be well described by the single-mode approximation.
@@ II.1 Quadrature squeezing and atomic homodyne detection
For the short-time evolution starting from polar state, atom number in the $|F=1,m=0\rangle$ component is approximately undepleted. At $q=|c_{2}|$, the complete Hamiltonian approximately takes the two-mode squeezing form after the replacement of $\hat{a}_{0}^{\dagger}\rightarrow\sqrt{N}e^{i\phi_{0}}$
@@
with $\phi_{0}=0$. Equation (S7) thus gives the equivalent two-mode squeezing dynamics as in undepleted pump optical four-wave mixing (FWM) or parametric down conversion, and the relevant angular momentum operators $L_{x}$ and $L_{y}$ as well as the nematic tensor operators $Q_{yz}$ and $Q_{xz}$ all reduce to quadrature operators
@@
In the simultaneous measurement scheme adopted for our experiment as depicted in Fig. 1(b) of the main text, the operations in the red dashed box execute the transformation of
@@
Here for simplicity, we take $\varepsilon_{1}\approx\varepsilon_{2}\approx 1/\sqrt{2}$, while in realistic experiments a small number of atoms (less than 100010001000) could remain in the $|F=1,m=0\rangle$ component, which can be neglected as the population of the pump is relatively large. Analogous to optical homodyne detection, $a_{1,\pm 1}$, or equivalently $a_{S(A)}$, is the signal field that contains the squeezing. While $a_{2,\pm 1}$ is the coherent pump, often referred to as the “local oscillator” in the context of optical homodyne detection.
The next two step coupling the signals $a_{S}$ and $a_{A}$ to the coherent pump enables atomic homodyne detection simultaneously. After the microwave (MW) pulses are applied, we obtain
@@
II.2 Experimental MW pulse sequence
Figure S1: The actual MW pulse sequence for simultaneous measurement of $Q_{yz}$ and $Q_{xz}$.
Limited by our experimental setup, the actual MW pulses applied for simultaneous measurement of $Q_{yz}$ and $Q_{xz}$ experimentally is shown in Fig. S1. The lower panel shows the MW pulse measurement sequence and the upper panel illustrates atom number distributions in all sublevels before (empty shape) and after (filled shape) the MW pulse sequence is implemented. The complete measurement protocol lasts for about 374 $\mu$s over a single experiment run. The blue, yellow, and green squares correspond to MW pulses with $\sigma_{+}$, $\sigma_{-}$, and $\pi-$polarizations, respectively. The blue and red shapes denote the signals and pump atomic modes for atomic homodyne detection, while the purple shapes depict the atoms after $\pi/2$-pulse couplings between the signals and the pumps. A small number of atoms remain in the $|F=1,m_{F}=0\rangle$ state as a result of imperfect pulse sequence. Here $C^{i,j}_{\varPhi}$ denotes the coherent coupling between $\left|F=1,m_{F}=i\right\rangle$ and $\left|F=2,m_{F}=j\right\rangle$ at a rotation angle $\varPhi$. For the first three pulses, $\varPhi_{1}$,$\varPhi_{2}$, and $\varPhi_{3}$ are adjusted by changing the power ratio between $\sigma_{+}$ and $\sigma_{-}$ polarizations to approximately split the atoms in $\left|F=1,m_{F}=0\right\rangle$ into two equal halves coherently.
Next, we describe the pulses in detail below. The $\pi/2$ coherent coupling between (1, ±1) and the coherent splitting of (1, 0) into (2, ±1) are achieved by the first three microwave pulses with only a small number of atoms (less than 1000) remaining in (1,0), using $F=2$ manifold as ancillary.
  1. 1.
    The first pulse facilitates a $\pi$-pulse coupling between (1, -1) and (2, 0), transferring atoms from (1, -1) to (2,0). Given that the microwave resonant with the transition between (1, -1) and (2,0) is also in resonance with the transition between (1, 0) and (2, -1), albeit with different polarizations, a part of atoms in (1, 0) is concurrently transferred to state (2, -1) as well.
  2. 2.
    The second microwave pulse induces a $\pi/2$-pulse coupling between (2, 0) and (1, 1), which forms $a_{S}=(a_{1}+a_{-1})/\sqrt{2}$ and $a_{A}=(a_{1}-a_{-1})/\sqrt{2}$. Since the microwave frequency resonant with the transition from (2, 0) to (1, 1) also matches the resonance condition for the transition between (1, 0) and (2, 1), a part of the atomic population in (1, 0) is transferred to (2, 1) as a result of this pulse.
  3. 3.
    The third microwave pulse replicates the $\pi$-pulse coupling between (1, -1) and (2, 0), and also transfers a part of the atoms in (1, 0) to (2, -1).
  4. 4.
    The fourth and fifth microwave pulses induce $\pi/2$-pulse couplings between (2, ±1) and (1, ±1), which enable the homodyne couplings between $a_{S(A)}$ and the coherent pump.
  5. 5.
    The sixth and seventh microwave pulses facilitate $\pi$-pulse couplings between (1, ±1) and (2, ±2), transferring atoms from (1, ±1) to (2, ±2) for subsequent absorption imaging.
The above microwave pulses are synthesized by signals radiated from two different antennas, whose $\sigma_{+/-}$ polarization ratios are different from each other. By tuning the relative phase and the amplitude of the two signals, we can adjust the polarization of the synthesized pulse. However, it is important to note that our microwave generators operate optimally within a restricted power range, which complicates the generation of pulses with a pure $\sigma_{+}$ or $\sigma_{-}$ polarization at the desired amplitude.
Figure S2: The numerical simulation of squeezing parameter measured simultaneously in both subspaces (without considering technical noises). The black dashed line is the inherent squeezing parameter of the spin-nematic squeezed state. The blue and red triangles correspond to the numerical simulation of our method for measuring squeezing in both subspaces. The blue and red circles correspond to the numerical simulation using the method in Ref. [49]. The parameters used for numerical simulation are identical to those in our manuscript.
Similar to the method in Ref. [49], our measurements utilize the ancillary $F=2$. However, our implemented measurement pulse sequence and the results we obtain significantly differ from the earlier work. The method reported in Ref. [49] is restricted to achieving detectable squeezing of up to 3 dB beyond the standard quantum limit (SQL). The partial degradation of detected squeezing is attributed to the transfer of half of the atoms from in $|F=1,m_{F}=\pm 1\rangle$ to the initially empty $|F=2,m_{F}=\pm 1\rangle$. Such a process mixes the squeezed mode $a_{1,\pm 1}$ with the vacuum mode, consequently amplifying the quantum noise of the squeezed mode. In contrast, our developed measurement sequence is theoretically capable of detecting quantum squeezing close to the quantum state’s intrinsic variance of the observables, thereby achieving near-optimal sensitivities. This enhanced performance is due to the ability of our designed sequence, as depicted in Fig. S1, to effectively prevent the mixing of squeezing modes with the vacuum noise. In Fig. S2, we show the numerical simulations of measuring spin-nematic squeezing in both SU(2) subspaces using the two methods, without considering other technical noises. The comparison between the two methods is obvious. The black dashed line represents the squeezing parameter of the quantum state as system evolves over time. The blue and red triangles illustrate the numerical simulation of our method for measuring squeezing in both subspaces, which approximates the squeezing of the spin-nematic squeezed state generated in our system. The blue and red circles correspond to the numerical simulation of the squeezing parameter using method in Ref. [49], demonstrating that the maximum achievable squeezing is limited to 3 dB.
II.3 Influence of crosstalk between “pump” and “signal”